A​ person's website specializes in the sale of rare or unusual vegetable seeds. He sells packets of​ sweet-pepper seeds for ​$2.16 each and packets of​ hot-pepper seeds for ​$4.40 each. He also offers a 16​-packet mixed pepper assortment combining packets of both types of seeds at ​$3.14 per packet. How many packets of each type of seed are in the​ assortment?
Answer by Boreal(15235)

S=x
H=16-x
2.16*x+4.40(16-x)=16(3.14)=50.24
2.16x+70.4-4.4x=50.24
-2.24x=-20.16
x=9 packets of sweet pepper
16-x=7 packets of hot-pepper